OpenStudy (anonymous):

An executive invests $24,000, some at 6% and some at 4% annual interest. If he receives an annual return of $1320, how much is invested at each rate?

OpenStudy (gw2011):

Let x=Amount invested at 6% Let y=Amount invested at 4% x+y=24,000 0.06x+0.04y=1320 Solve these equations by multiplying the top one by 0.06 to eliminate x and find y: 0.06x+0.06y=1440 0.06x+0.04y=1320 Subtract this equation from the top one to get: 0.02y=120 y=$6000 invested at 4% Substitute 6000 into x+y=24,000 to find x: x+y=24,000 x+6000=24,000 x=24,000-6000 x=$18,000 invested at 6% Therefore, $6000 is invested at 4% and $18,000 is invested at 6%